OpenStudy (danmac0710):

i used proof by induction to prove the following but I think I may have missed an easier method... would you mind having a look please?

OpenStudy (thomas5267):

Question?

OpenStudy (danmac0710):

sorry i am editing it

OpenStudy (danmac0710):

\[\frac{ 6^n }{ n! }\le \frac{ 6^5 }{ 5! }\times \frac{ 6 }{ n }\]

OpenStudy (danmac0710):

for n greater than 6

OpenStudy (danmac0710):

forgive my slow editing. new computer and tablet and struggling a bit!

OpenStudy (ikram002p):

its clear by induction

OpenStudy (danmac0710):

that was how i did it, but i wondered if I mised an easier way. Thanks!

OpenStudy (zzr0ck3r):

the best thing to do for proving something is true for all n, is induction.